OpenStudy (anonymous):

Refresh my memory: What is the formula for the difference of two squares? Two cubes?

OpenStudy (hoblos):

a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

Yes, and the difference of two cubes??

OpenStudy (hoblos):

a^3 - b^3 = (a-b)(a^2 +ab + b^2)

OpenStudy (anonymous):

Thanks! @hoblos

OpenStudy (hoblos):

yw :)